OpenStudy (anonymous):

A square window has an area of 96 square inches. What is the length of each side of the window. Help

OpenStudy (anonymous):

What is the equation for the area of a square?

OpenStudy (anonymous):

side*side right?

OpenStudy (anonymous):

so if the area of the square is 96 than side*side = 96; now solve for side

OpenStudy (anonymous):

yea but it has to be in simplest radical form

OpenStudy (anonymous):

ah okay so sqrt(96) 96 = 12*8 = (2*2*3)*(2*2*2) right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

k so sqrt(2*2*2*2*2*3) = sqrt[(2*2)*(2*2)*2*3] = (2)*(2)*sqrt(2*3) understand how I wrote this?

OpenStudy (anonymous):

so answer is \[4*\sqrt{6}\]

OpenStudy (anonymous):

ok i get it thank you so much

OpenStudy (anonymous):

your welcome, all you are looking for is pairs of multiples of the number under the radical

OpenStudy (anonymous):

Thank you ill remember that